OpenStudy (anonymous):

What is the solution to g over two is equal to one over five? 0.1 0.4 4 10

OpenStudy (anonymous):

g/2=1/5 (g/2)*2=(1/5)*2 g=2/5 2/5=0.4 Confirmation: http://www.wolframalpha.com/input/?i=g%2F2%3D1%2F5&dataset= :)

OpenStudy (anonymous):

thank you so much:)

OpenStudy (anonymous):

No problem :)